# End-to-End Conformal Calibration for Optimization Under Uncertainty

### v1

@@ Abstract
Machine learning can significantly improve performance for decision-making under uncertainty in a wide range of domains. However, ensuring robustness guarantees requires well-calibrated uncertainty estimates, which can be difficult to achieve in high-capacity prediction models such as deep neural networks. Moreover, in high-dimensional settings, there may be many valid uncertainty estimates, each with their own performance profile—i.e., not all uncertainty is equally valuable for downstream decision-making. To address this problem, this paper develops an end-to-end framework to learn the uncertainty estimates for conditional robust optimization, with robustness and calibration guarantees provided by conformal prediction. In addition, we propose to represent arbitrary convex uncertainty sets with partially input-convex neural networks, which are learned as part of our framework. Our approach consistently improves upon two-stage estimate-then-optimize baselines on concrete applications in energy storage arbitrage and portfolio optimization.
@@ 1 Introduction
Well-calibrated estimates of forecast uncertainty are vital for risk-aware decision-making in many real-world systems. For instance, grid-scale battery operators forecast electricity prices to schedule battery charging/discharging to maximize profit, while they rely on uncertainty estimates to minimize financial or operational risk. Similarly, financial investors use forecasts of asset returns along with uncertainty estimates to maximize portfolio returns while minimizing downside risk.
Historically, approaches for decision-making under uncertainty have often treated the estimation of uncertainty separately from its use for downstream decision-making. This “estimate then optimize” (ETO) paradigm [13] separates the problem into an “estimate” stage, where a predictive model is trained to forecast the uncertain quantity, yielding an uncertainty set estimate, followed by an “optimize” stage, where the forecast uncertainty is used to make a decision. Notably, any cost or loss associated with the downstream decision is usually not provided as feedback to the predictive model.
A recent line of work [13, 32, 42, 12] has made steps toward bridging the gap between uncertainty quantification and robust optimization-driven decision-making, where optimization problems take a forecast uncertainty set as a parameter, as is common in energy systems [45, 31] and financial applications [21, 34]. However, existing approaches are suboptimal for several reasons:
@@
    The predictive model is not trained with feedback from the downstream objective. Because the downstream objective is often asymmetric with respect to the forecasting model’s error, the trained predictive model, though it provides accurate predictions, may yield significantly worse performance on the true decision-making objective.
@@
    For the robust optimization to be tractable, the forecast uncertainty sets have restricted parametric forms. Common parametric forms include box and ellipsoidal uncertainty sets, limiting the expressivity of uncertainty estimates.
@@
    Because neural network models are often poor at estimating their own uncertainty, the forecasts may not be well-calibrated. Recent approaches such as isotonic regression [24] and conformal prediction [38] have made progress in providing calibrated uncertainty estimates from deep learning models, but such methods are typically applied post-hoc to pre-trained models and are therefore difficult to incorporate into an end-to-end training procedure.
@@
    We develop a methodology for training prediction models end-to-end with downstream decision-making objectives and conformally calibrated uncertainty estimates in the context of the conditional robust optimization problem. This framework is illustrated in Figure 1. By including differentiable conformal calibration layers in our model during training, we close the loop and ensure that feedback from the uncertainty’s impact on the downstream objective is accounted for in the training process, since not all model errors nor uncertainty estimates will result in the same downstream cost. This end-to-end training enables the model to focus its learning capacity on minimizing error and uncertainty on outputs with the largest decision-making cost, with more leeway for outputs that have lower costs.
@@
    We propose using partially input-convex neural networks (PICNNs) [4] to approximate arbitrary convex uncertainty sets for the conditional robust optimization problem. Due to the universal convex function approximation property these networks enjoy [11], this approach enables training far more general representations of uncertainty than prior works have considered, which in turn yields substantial improvements on downstream decision-making performance. Importantly, PICNNs are well-matched to our conditional robust optimization problem, as we show that the robust problem resulting from this parametrization can be tractably reformulated as a convex optimization problem.
@@
    We propose an exact and computationally efficient method for differentiating through the conformal prediction procedure during training. In contrast to prior work [40], our method gives exact gradients, without relying on approximate ranking and sorting methods.
Finally, we perform extensive experiments evaluating the performance of our approach on two applications: an energy storage arbitrage task and a portfolio optimization problem. Our code is available on GitHub. [^1] https://github.com/chrisyeh96/e2e-conformal We demonstrate conclusively that the combination of end-to-end training with the flexibility of the PICNN-based uncertainty sets consistently improves over ETO baseline methods.
Figure 1: Our proposed framework for end-to-end conformal calibration for optimization under uncertainty updates the machine learning model using gradients from the task loss.
@@ 2 Problem Statement and Background
Our problem setting is defined formally as follows: suppose that data $(x,y)\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ is sampled i.i.d. from an unknown joint distribution $\mathcal{P}$. Upon observing the input $x$ (but not the label $y$), an agent makes a decision $z\in\mathbb{R}^{p}$. After the decision is made, the true label $y$ is revealed, and the agent incurs a task loss $f(x,y,z)$, for some known task loss function $f:\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{p}\to\mathbb{R}$. In addition, the agent’s decision must satisfy a set of joint constraints $g(x,y,z)\leq 0$ coupling $x,y$, and $z$.
As an illustrative example, consider an agent who would like to minimize the costs of charging and discharging a battery. The agent may use weather forecasts and historical observations $x$ to predict future energy prices $y$. Based on the predicted prices, the agent decides on the amount $z$ to charge or discharge the battery. The task loss $f$ is the cost incurred by the agent, and the constraints $g$ include limits on how fast the battery can charge as well as the maximum capacity of the battery. This example is explored in more detail in Section 5.
Because the agent does not observe the label $y$ prior to making its decision, ensuring good performance and constraint satisfaction requires that the agent makes decisions $z$ that are robust to the various outcomes of $y$. A common objective is to choose $z$ to robustly minimize the task loss and satisfy the constraints over all realizations of $y$ within a $(1-\alpha)$-confidence region $\Omega(x)\subset\mathbb{R}^{n}$ of the true conditional distribution $\mathcal{P}(y\mid x)$, where $\alpha\in(0,1)$ is a fixed risk level chosen based on operational requirements. In this case, the agent’s robust decision can be expressed as the optimal solution to the following conditional robust optimization (CRO) problem [13, 12]:
$z^{\star}(x):=\operatorname*{arg\,min}_{z\in\mathbb{R}^{p}}\max_{\hat{y}\in\Omega(x)}f(x,\hat{y},z)\quad\text{s.t. }g(x,\hat{y},z)\leq 0.$ (1)
After the agent decides $z^{\star}(x)$, the true label $y$ is revealed, and the agent incurs the task loss $f(x,y,z^{\star}(x))$. Thus, the agent seeks to minimize the expected task loss
@@
While the joint distribution $\mathcal{P}$ is unknown, we assume that we have access to a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ of samples from $\mathcal{P}$. Then, our objective is to train a machine learning model to learn an approximate $(1-\alpha)$-confidence set $\Omega(x)$ of possible $y$ values for each input $x$. Formally, our learned $\Omega(x)$ should satisfy the following marginal coverage guarantee.
Definition 1 (marginal coverage).
@@
[…]
Closest to our current work are recent papers [12, 42] which proposed “end-to-end” (E2E) formulations of the CRO problem. Specifically, Chenreddy and Delage [12] poses a training methodology for learning conditional uncertainty sets $\Omega(x)$ for CRO that incorporates a “conditional coverage loss” into the downstream objective, thereby promoting calibrated uncertainty while training $\Omega(x)$ based on the task loss. On the other hand, Wang et al. [42] approaches the end-to-end CRO problem by reformulating the inner maximization problem in their version of (1) into a relaxed form using the Conditional Value-at-Risk (CVaR) and solving for a single uncertainty set $\Omega$ that is independent of $x$. While these papers take significant steps toward learning uncertainty sets end-to-end for conditional robust optimization, they both have substantial limitations. Both papers consider restrictive parametric forms of uncertainty, limiting the performance improvement that their methods can achieve: Chenreddy and Delage [12] focus solely on ellipsoidal uncertainty sets, and Wang et al. [42] restrict to canonical forms of uncertainty sets from the robust optimization literature such as box, ellipsoidal, and polyhedral uncertainty. In addition, the latter work does not learn $x$-dependent uncertainty sets, limiting their ability to adapt to heteroscedasticity in $\mathcal{P}(y\mid x)$ and yielding conservative behavior. Moreover, the conditional coverage loss approach in Chenreddy and Delage [12] does not ensure provable coverage guarantees for their learned uncertainty sets, while the CVaR relaxation in Wang et al. [42] will typically yield additional conservativeness.
@@
In contrast, our work avoids these limitations to obtain a more optimal and flexible methodology: we incorporate differentiable conformal calibration layers into our training procedure along with a standard split conformal post-hoc calibration step during inference, which provide provable guarantees on coverage while ensuring that uncertainty is learned end-to-end in a manner that is both calibrated and specifically focused on minimizing the task loss. Furthermore, we propose to use partially input-convex neural networks (PICNNs) [4], which provide a general, expressive, and nonparametric representation of conditional convex uncertainty regions that can vary with $x$ and be used efficiently in robust optimization.
@@
Beyond the above closely related work, this paper builds upon and contributes to several different areas in machine learning and robust optimization; see Appendix A for a comprehensive discussion.
@@ 3 End-to-End Training of Conformally Calibrated Uncertainty Sets
In this section, we describe our proposed methodological framework for end-to-end task-aware training of predictive models with conformally calibrated uncertainty for the conditional robust optimization problem (1). Figure 1 illustrates our framework, and Algorithm 1 shows pseudocode for both training and inference. Our overarching goal is to learn uncertainty sets $\Omega(x)$ which provide $(1-\alpha)$ coverage for any choice of $\alpha\in(0,1)$, and which offer the lowest possible task loss (2).
To this end, we must consider three primary questions:
@@
In this section, we will answer the latter two questions when faced with common parametric forms of the uncertainty set $\Omega(x)$ often used in the literature. Then, in Section 4, we will describe our proposed approach of using PICNNs to represent general convex uncertainty sets.
@@
Throughout the rest of the paper, we make the following assumptions on the functions $f$ and $g$ and the uncertainty set $\Omega$ in (1) to ensure tractability of the resulting optimization problem.
Assumptions.
We assume that the constraint $g(x,y,z)=g(x,z)$ does not depend on $y$, and that the task loss is of the form $f(x,y,z)=y^{\top}Fz+\tilde{f}(x,z)$ for some matrix $F\in\mathbb{R}^{n\times p}$ and auxiliary function $\tilde{f}:\mathbb{R}^{m}\times\mathbb{R}^{p}\to\mathbb{R}$—that is, the task loss decomposes into a function $\tilde{f}(x,z)$ independent of $y$ and a bilinear term $y^{\top}Fz$. We further assume that the functions $g(x,z)$ and $\tilde{f}(x,z)$ are convex in $z$, and that for each $x\in\mathbb{R}^{m}$, the uncertainty set $\Omega(x)$ is a convex set.
Note that some of these assumptions can be relaxed under various conditions on the uncertainty set $\Omega(x)$ while maintaining tractability of (1). For example, if $\Omega(x)$ is expressed as the convex hull of a finite set of vertices, then $g(x,y,z)$ can represent a conic constraint (see, e.g., [6, Chapter 6]).

### v2

@@ Abstract
Machine learning can significantly improve performance for decision-making under uncertainty across a wide range of domains. However, ensuring robustness guarantees requires well-calibrated uncertainty estimates, which can be difficult to achieve with neural networks. Moreover, in high-dimensional settings, there may be many valid uncertainty estimates, each with its own performance profile—i.e., not all uncertainty is equally valuable for downstream decision-making. To address this problem, this paper develops an end-to-end framework to learn uncertainty sets for conditional robust optimization in a way that is informed by the downstream decision-making loss, with robustness and calibration guarantees provided by conformal prediction. In addition, we propose to represent general convex uncertainty sets with partially input-convex neural networks, which are learned as part of our framework. Our approach consistently improves upon two-stage estimate-then-optimize baselines on concrete applications in energy storage arbitrage and portfolio optimization.
@@ 1 Introduction
Well-calibrated estimates of forecast uncertainty are vital for risk-aware decision-making in many real-world systems. For instance, grid-scale battery operators forecast electricity prices to schedule battery charging/discharging to maximize profit, while accounting for forecast uncertainty to mitigate financial or operational risk. Similarly, financial investors use forecasts of asset returns with uncertainty estimates to maximize portfolio returns while minimizing downside risk.
Traditional approaches to decision-making under uncertainty often follow an “estimate-then-optimize” (ETO) paradigm (chenreddy_data-driven_2022), in which the uncertainty estimation and decision-making stages are decoupled. In the “estimation” stage, a predictive model is trained to forecast a target quantity along with an uncertainty set. In the “optimization” stage, this uncertainty estimate informs a downstream decision. Crucially, the cost or performance of the downstream decision is typically not fed back into the training of the predictive model.
A recent line of work (chenreddy_data-driven_2022; patel_conformal_2024; wang_learning_2024; chenreddy_end--end_2024) has made steps toward bridging the gap between uncertainty quantification and robust optimization-driven decision-making, where optimization problems take a forecast uncertainty set as a parameter, as is common in energy systems (yan_robust_2020; parvar_optimal_2022) and financial applications (gregory_robust_2011; quaranta_robust_2008). However, existing approaches are suboptimal for several reasons:
@@
    Lack of decision-aware training: The predictive model is typically not trained with feedback from the downstream objective. Because the downstream objective is often asymmetric with respect to the forecasting model’s errors, prediction models trained with decision-agnostic losses may achieve high predictive accuracy but still perform poorly on the decision-making objective.
@@
    Restricted uncertainty set parameterizations: To preserve tractability of the robust optimization problem, uncertainty sets are often constrained to simple parametric forms (e.g., boxes or ellipsoids), limiting the expressivity of uncertainty estimates.
@@
    Calibration challenges: Because neural network models are often poor at estimating their own uncertainty, the forecasts may not be well-calibrated. Recent approaches such as isotonic regression (kuleshov_accurate_2018) and conformal prediction (shafer_tutorial_2008) have made progress in providing calibrated uncertainty estimates from deep learning models, but such methods are typically applied post-hoc to trained models and are therefore difficult to incorporate into an end-to-end training procedure.
Figure 1: Whereas prior “estimate-then-optimize” (ETO, top) approaches separate the model training from the optimization (decision-making) procedure, we propose a framework for end-to-end (E2E, bottom) conformal calibration for optimization under uncertainty that directly trains the machine learning model using gradients from the task loss.
@@
    We develop a framework for training prediction models end-to-end with downstream decision-making objectives and conformal-calibrated uncertainty sets in the context of the conditional robust optimization problem. This framework is illustrated in Figure˜1 (bottom). By including differentiable conformal calibration in our model during training, we close the loop and ensure that feedback from the uncertainty’s impact on the downstream objective is accounted for in the training process, since not all model errors nor uncertainty estimates will result in the same downstream cost. This end-to-end training enables the model to focus its learning capacity on minimizing error and uncertainty on outputs with the largest decision-making cost, with more leeway for outputs that have lower costs.
@@
    We propose using partially input-convex neural networks (PICNNs) as the nonconformity score function for conformal prediction, enabling the approximate parametrization of arbitrary compact, convex uncertainty sets in the conditional robust optimization problem. To the best of our knowledge, no existing works use PICNNs to parametrize such arbitrary convex uncertainty sets. Due to the universal convex function approximation property these networks enjoy (chen_optimal_2019), this approach enables training much more general representations of uncertainty than prior works have considered, which in turn yields substantial improvements on downstream decision-making performance. Importantly, PICNNs are well-matched to our conditional robust optimization problem: we show that the robust problem resulting from this parametrization can be reformulated as a tractable convex optimization problem.
@@
    We propose an exact and computationally efficient method to differentiate through the conformal prediction procedure during training. Unlike prior work (stutz_learning_2022), our method gives exact gradients, without using approximate ranking and sorting methods.
Finally, we evaluate the performance of our approach on two applications: an energy storage arbitrage task and a portfolio optimization problem. We demonstrate that the combination of end-to-end training with the flexibility of the PICNN-based uncertainty sets consistently improves over ETO baseline methods. The performance benefit of our end-to-end method is apparent even under distribution shift.
@@ 2 Problem Statement and Background
Our problem is defined formally as follows: suppose that data $(x,y)\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ is sampled i.i.d. from an unknown joint distribution $\mathcal{P}$. Upon observing the input $x$ (but not the label $y$), an agent makes a decision $z\in\mathbb{R}^{p}$. After the decision is made, the true label $y$ is revealed, and the agent incurs a task loss $f(x,y,z)$, for some known task loss function $f:\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{p}\to\mathbb{R}$. In addition, the agent’s decision must satisfy a set of joint constraints $g(x,z)\leq 0$ coupling $x$ and $z$.
As an illustrative example, consider an agent who would like to minimize the costs of charging and discharging a battery over 24 hours in a day. The agent may use weather forecasts and historical observations $x$ to predict future energy prices $y$. Based on the predicted prices, the agent decides on the amount $z$ to charge or discharge the battery. The task loss $f$ is the cost incurred by the agent, and the constraints $g$ include limits on how fast the battery can charge as well as the maximum capacity of the battery. This example is explored in more detail in Section˜5.
Because the agent does not observe the label $y$ prior to making its decision, ensuring good performance and constraint satisfaction requires that the agent makes decisions $z$ that are robust to the various outcomes of $y$. A common objective is to choose $z$ to robustly minimize the task loss and satisfy the constraints over all realizations of $y$ within a $(1-\alpha)$-confidence region $\Omega(x)\subset\mathbb{R}^{n}$ of the true conditional distribution $\mathcal{P}(y\mid x)$, where $\alpha\in(0,1)$ is a fixed risk level based on operational requirements. In this case, the agent’s robust decision can be expressed as the optimal solution to the following conditional robust optimization (CRO) problem (chenreddy_data-driven_2022):
$\hskip-6.00006ptz^{\star}(x):=\operatorname*{arg\,min}_{z\in\mathbb{R}^{p}}\max_{\hat{y}\in\Omega(x)}f(x,\hat{y},z)\ \text{s.t. }g(x,z)\leq 0.$ (1)
After the agent decides $z^{\star}(x)$, the true label $y$ is revealed, and the agent incurs the task loss $f(x,y,z^{\star}(x))$. Thus, the agent seeks to minimize expected task loss
@@
While the joint distribution $\mathcal{P}$ is unknown, we assume that we have a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ of i.i.d. samples from $\mathcal{P}$. Then, our objective is to train a machine learning model to learn an approximate $(1-\alpha)$-confidence set $\Omega(x)$ of possible $y$ values for each input $x$. Formally, our learned $\Omega(x)$ should satisfy the following marginal coverage guarantee.
Definition 2.1 (marginal coverage).
@@
The problem of constructing data-driven and machine-learned uncertainty sets with probabilistic coverage guarantees for use in robust optimization has been widely explored in prior literature (e.g., bertsimas_robust_2006; bertsimas_data-driven_2018; alexeenko_nonparametric_2020; goerigk_data-driven_2023). chenreddy_data-driven_2022 first coined the phrase “conditional robust optimization” for the problem (1) and considered learning context-dependent uncertainty sets $\Omega(x)$ in this setting. However, their approach results in a mixed integer optimization that is intractable to solve for large-scale problems. Moreover, they follow the “estimate then optimize” (ETO) paradigm (elmachtoub_estimate-then-optimize_2023). As shown in Figure˜1 (top), the ETO paradigm separates the machine learning model training from the decision optimization. The lack of feedback from the downstream task loss during model training in ETO generally leads to uncertainty sets $\Omega(x)$ which yield suboptimal results. Several other recent papers follow the ETO paradigm using homoskedastic ellipsoidal uncertainty sets (johnstone_conformal_2021), heteroskedastic box and ellipsoidal uncertainty sets (sun_predict-then-calibrate_2023), and a “union of balls” parametrization of uncertainty (patel_conformal_2024). In our experiments (Section˜5), we demonstrate consistent improvements over the methods of johnstone_conformal_2021 and sun_predict-then-calibrate_2023.
Closest to our work is an “end-to-end” formulation of the CRO problem posed by chenreddy_end--end_2024, which aims to learn conditional uncertainty sets $\Omega(x)$ using a weighted combination of the downstream task loss along with a “conditional coverage loss” to promote calibrated uncertainty. However, they focus solely on ellipsoidal uncertainty sets, and their conditional coverage loss does not provably ensure coverage for their learned uncertainty sets. In our experiments, we do not compare against chenreddy_end--end_2024 because we only consider other methods with a provable coverage guarantee. [^1] As of the time of writing, the approach of chenreddy_end--end_2024 also suffers from substantial inconsistencies between their code implementation and the equations from their paper. In particular, the conditional coverage loss proposed in their paper is not implementable, as it will (almost surely) yield zero gradients.
A concurrent related work by wang_learning_2024 approaches the problem of learning unconditional uncertainty sets for robust optimization, while achieving robust constraint satisfaction guarantees. While their method also uses an end-to-end task loss, we find their use of the same uncertainty set $\Omega$ for every problem instance (i.e., $\Omega$ is independent of $x$) to be highly restrictive. For example, in the context of our battery control problem, this restriction would disallow the use of weather forecasts and historical price data to estimate uncertainty in future energy prices. [^2] After the present paper’s submission, the authors of wang_learning_2024 updated their preprint and expanded their formulation to accommodate conditional uncertainty sets. They also use restrictive uncertainty set parametrizations such as box, ellipsoidal, and polyhedral uncertainty. Moreover, their guarantee on constraint satisfaction requires an assumption on the convergence of their algorithm to an approximate stationary point; however, this convergence is only guaranteed asymptotically, and not in finitely many samples.
An alternative line of research designs post-hoc conformal prediction sets aimed at minimizing decision costs in robust classification problems (kiyani_decision_2025; cortes-gomez_utility-directed_2024). However, to date, it is unclear how to adapt these methods, which produce discrete uncertainty sets, for regression settings.
@@
In contrast, our work overcomes these limitations: we incorporate differentiable conformal calibration during training to ensure that uncertainty is learned end-to-end in a manner that is both calibrated and minimizes task loss. We apply split conformal post-hoc calibration during inference for provable guarantees on coverage. Furthermore, we use partially input-convex neural networks (amos_input_2017) to directly parameterize the nonconformity score function in conformal prediction, enabling a general and expressive representation of arbitrary conditional convex uncertainty regions that can vary with $x$ and be used efficiently in robust optimization.
@@
Beyond the above closely related work, this paper builds upon and contributes to several different areas in machine learning and robust optimization; see Appendix˜B for a comprehensive discussion.
@@ 3 End-to-End Training of Conformally Calibrated Uncertainty Sets
In this section, we describe our proposed methodological framework for end-to-end task-aware training of predictive models with conformally calibrated uncertainty for the conditional robust optimization problem (1). Our overarching goal is to learn uncertainty sets $\Omega(x)$ which provide $(1-\alpha)$ coverage for any choice of $\alpha\in(0,1)$, and which offer the lowest possible task loss (2). To this end, we must consider three primary questions:
@@
Figure˜1 (bottom) illustrates the key parts of our framework to answer these questions. First, we use a machine learning model to parametrize a nonconformity score function $s_{\theta}$, and we define the uncertainty set $\Omega(x)$ to be a $q$-sublevel set of $s_{\theta}(x,\cdot)$. Second, we use conformal calibration to pick $q$ to enforce marginal coverage. Third, we backpropagate gradients through both the robust optimization and conformal calibration steps to update the machine learning model, thereby enabling end-to-end learning. Sections˜3.1, 3.2 and 3.3 describe each of these parts in detail, and Algorithm˜1 shows pseudocode for both training and inference.
@@
For the rest of the paper, we make the following assumptions on the functions $f$ and $g$ to ensure tractability of the resulting optimization problem.
Assumption 3.1.
We assume the task loss has the form $f(x,y,z)=y^{\top}F(x,z)+\tilde{f}(x,z)$, where $F(x,z)$ is an affine function of $z$ and $\tilde{f}(x,z)$ is convex in $z$. Furthermore, we assume that $g(x,z)$ is convex in $z$.
Technically, the assumption on $f$ can be relaxed to the more general setting where $f$ is a conic-representable saddle function that is convex in $z$ and concave in $y$ (juditsky_well-structured_2022), which admits an automated dual representation (schiele_disciplined_2023). However, we believe the core ideas for our exposition and proofs are much more straightforward with the stronger assumption in ˜3.1.
3.1 Representations of the uncertainty set
We consider convex uncertainty sets of the form
[…]
In choosing a particular score function $s_{\theta}$, one must balance two considerations: first, the generality of the sets $\Omega_{\theta}(x)$ that $s_{\theta}$ can represent, and second, the tractability of the resulting robust optimization problem (1). We will now show that our representation (3) generalizes commonly-used box and ellipsoidal uncertainty sets, which are known to have tractable robust problems; later, in Section˜4, we will propose to approximate more general convex uncertainty sets using partially input-convex neural networks.
Box uncertainty sets.
A simple uncertainty representation is box uncertainty where $\Omega(x)=[\underline{y}(x),\overline{y}(x)]$ is an $n$-dimensional box whose lower and upper bounds depend on $x$. Let $h_{\theta}:\mathbb{R}^{m}\to\mathbb{R}^{n}\times\mathbb{R}^{n}$ be a neural network that estimates lower and upper bounds: $h_{\theta}(x)=(h_{\theta}^{\text{lo}}(x),h_{\theta}^{\text{hi}}(x))$. To represent a box uncertainty set in the form (3), we define a nonconformity score function that generalizes scalar conformalized quantile regression (romano_conformalized_2019):
$s_{\theta}(x,y)=\max(\left\lVert h_{\theta}^{\text{lo}}(x)-y\right\rVert_{\infty},\left\lVert y-h_{\theta}^{\text{hi}}(x)\right\rVert_{\infty}).$
Then, the uncertainty set (3) becomes
$\Omega_{\theta}(x)=\left[h_{\theta}^{\text{lo}}(x)-q\mathbf{1},\ h_{\theta}^{\text{hi}}(x)+q\mathbf{1}\right]=:\Big[\underline{y}(x),\ \overline{y}(x)\Big].$
Given a box uncertainty set $\Omega_{\theta}(x)$, we can take the dual of the inner maximization problem (see Section˜C.1) to transform the robust optimization problem (1) into an equivalent form that is convex, and hence tractable, under ˜3.1:
$z_{\theta}^{\star}(x)=\operatorname*{arg\,min}_{z\in\mathbb{R}^{p}}\min_{\nu\in\mathbb{R}^{n}}$ $(\overline{y}(x)-\underline{y}(x))^{\top}\nu+\underline{y}(x)^{\top}F(x,z)+\tilde{f}(x,z)$ (4)
s.t. $\nu\geq\mathbf{0},\ \nu-F(x,z)\geq\mathbf{0},\ g(x,z)\leq 0.$
Ellipsoidal uncertainty sets.
Another common form of uncertainty set is ellipsoidal uncertainty. Suppose a neural network model $h_{\theta}:\mathbb{R}^{m}\to\mathbb{R}^{n}\times\mathbb{S}_{+}^{n}$ predicts mean and covariance parameters $h_{\theta}(x)=(\mu_{\theta}(x),\Sigma_{\theta}(x))$, so that $\hat{\mathcal{P}}(y\mid x;\theta)=\operatorname{\mathcal{N}}(y\mid\mu_{\theta}(x),\Sigma_{\theta}(x))$ denotes a predicted conditional density, where $\operatorname{\mathcal{N}}(\cdot\mid\mu,\Sigma)$ is the multivariate normal density function. In this case, we define the nonconformity score function based on the squared Mahalanobis distance (johnstone_conformal_2021; sun_predict-then-calibrate_2023)
which yields uncertainty sets (3) that are ellipsoidal:
$\Omega_{\theta}(x)=\{\hat{y}\mid(\hat{y}-\mu_{\theta}(x))^{\top}(\Sigma_{\theta}(x))^{-1}(\hat{y}-\mu_{\theta}(x))\leq q\}.$
Let $L_{\theta}(x)$ denote the unique lower-triangular Cholesky factor of $\Sigma_{\theta}(x)$ (i.e., $\Sigma_{\theta}(x)=L_{\theta}(x)L_{\theta}(x)^{\top}$). Taking the dual of the inner maximization problem and invoking strong duality (see Section˜C.2), we transform the robust optimization problem (1) into an equivalent form that is convex, and hence tractable, under ˜3.1:
$z_{\theta}^{\star}(x)=\operatorname*{arg\,min}_{z\in\mathbb{R}^{p}}$ $\sqrt{q}\|L_{\theta}(x)^{\top}F(x,z)\|_{2}+\mu_{\theta}(x)^{\top}F(x,z)+\tilde{f}(x,z)$ (5)
s.t. $g(x,z)\leq 0,$